OpenStudy (anonymous):

Which of the following equations is used to find the value of c? a²+b²=c² c²+b²=a² a²+c²=b² a²+b²+c²=1

OpenStudy (tkhunny):

For what? It's different for ellipses and hyperbolas.

OpenStudy (anonymous):

ellipses sorry

OpenStudy (tkhunny):

Go look back at the last problem. The answer is in there. You must remember from one problem to the next! :-)

OpenStudy (anonymous):

@tkhunny you said b2=a2−c2 but that is not an option here

OpenStudy (tkhunny):

You must also remember your algebra. Here's what I said: b^2 = a^2 - c^2 That is the same as this: b^2 + c^2 = a^2 That is the same as this b^2 - a^2 = -c^2 That is the same as this: -a^2 = -b^2 - c^2 That is the same as this: c^2 = a^2 - b^2 Are you seeing it, yet?

OpenStudy (anonymous):

oh okay so it is the same?

OpenStudy (anonymous):

a^2+c^2=b^2 is the answer in that case? @tkhunny

OpenStudy (tkhunny):

No, you jut changed one sign. That's no good. This we have established: b^2 = a^2 - c^2 Notice how all four chhoices are absent any negative signs. Rewrite the known expression, making sure there is nothgin engative when you are done. b^2 = a^2 - c^2 is hte same as b^2 + c^2 = a^2 Do we have that one?

OpenStudy (anonymous):

ohhh okay i see now, yes that is an option thank you so much. i have just one or two more problems i need help with.. i'll repost now can you please help? @tkhunny